### v1

@@ Abstract
This paper introduces the 193-bus synthetic Korean power grid (KPG 193), developed using open data sources to address recent challenges of the Korean power system. The KPG 193 test system serves as a valuable platform for decarbonization research, capturing Korea’s low renewable energy penetration, concentrated urban energy demand, and isolated grid structure. Clustering techniques were applied to preserve key system characteristics while maintaining computational tractability and representativeness. The system includes 193 buses, 123 generators, 407 transmission lines, and incorporates temporal weather datasets. Its feasibility was validated through Unit Commitment (UC), DC Optimal Power Flow (DCOPF) and AC Optimal Power Flow (ACOPF) simulations using 2022 demand and renewable generation data. This test system aims to provide a foundational framework for modeling and analyzing the Korean power grid.
@@ I Introduction
As of 2022, the Asia-Pacific region accounts for over 50% of global greenhouse gas emissions, with more than 80% of the projected global coal demand growth originating from this region [1]. In Korea, coal contributes 33% of the country’s electricity supply, while renewables account for less than 10% [2]. Given the region’s significant impact on global emissions, advancing open-source power grid modeling efforts in Korea is crucial to support decarbonization studies. Europe has made substantial progress in open-source power grid modeling. The SciGrid [3] and osmTGmod [4] projects modeled Germany’s transmission network using OpenStreetMap (OSM), a publicly available dataset [5]. Building on this, PyPSA-Eur [6] extended the scope to create an open model covering the entire ENTSO-E area. A key enabler in Europe is the public availability of geographical data for substations and transmission lines provided by ENTSO-E [7]. This openness allows for the validation of power grid models built on OSM data, enhancing their accuracy and reliability. In the United States, open test systems have also achieved significant development. The pioneering work at Texas A&M University (TAMU) [8] laid the foundation, later updated the system to meet 2016 standards with additional generators and high-voltage direct current (HVDC) transmission lines [9]. More recently, the TX-123BT network, which represents the Texas power grid and incorporates five years of meteorological data, has been made publicly available [10].
@@
While regions such as Europe and the United States have made significant advancements in open-source power grid modeling through extensive research collaboration, synthetic power grid systems remain underdeveloped in Korea. Previous studies have utilized a realistic 1,428-bus system [11] and a reduced 11-zone system [12]; however, only the 11-zone system is publicly available [13]. To address this gap, we present the synthetic Korean power grid 193-bus test system (KPG 193), developed using open data from the year 2022. Detailed information about the Korean power grid is not publicly disclosed and publicly accessible data is limited to electricity demand and renewable energy capacity at the municipal level. To overcome these limitations while ensuring computational tractability and accurate grid representation, we developed the test system by clustering regions based on the locations of regional offices of the utility company. The proposed KPG 193 test system captures recent challenges of Korea’s power grid, including concentrated electricity demand in metropolitan areas and the geographical distribution of coal and nuclear power plants along the coastline. The main contributions of this paper are threefold:
@@
  • Development of the KPG 193 test system: We provide synthetic transmission grid model for Korea based on open-source data, featuring 193 buses, 123 generators, and 407 transmission lines.
@@
  • Integration of renewables and weather datasets: The system includes spatial and temporal weather datasets used to generate renewable generation profile, enabling accurate simulations and facilitating advanced energy system studies.
@@ II-A1 Network Topology
Two types of line data were extracted from the OSM database using the Overpass API: power line for overhead lines and power cable for underground cables [17]. Each line element includes attributes such as voltage level and the number of parallel lines. AC line parameters were calculated using standard conductor coefficients from manufacturer catalogs, considering voltage level and conductor type [18, 19]. A summary of these line parameters is provided in Table II. The Korean power grid also features a high-voltage direct current (HVDC) transmission line connecting the Bukdangjin and Godeok substations with a capacity of 1,500 MVA [20]. Since this HVDC line is not included in the OSM database, it was manually added during data preparation.
@@
Overhead line 154 kV ACSR 410 mm×$\times 2$ 2 0.074 0.337 12.9 452
345 kV ACSR 480 mm×$\times 2$ 2 0.062 0.332 13.1 1,086
@@
Capacity [GW] 38.90 37.80 25.20 23.75 7.20 1.65 134.49
Capacity Share [%] 29% 28% 19% 18% 5% 1% 100%
@@ II-B3 Renewable Energy Profile Generation
Renewable energy profiles were generated for each energy source using different methods. The hydropower profile $U_{t}$ was derived from historical generation data provided by KPX [27], representing uniform hydroelectric production across all buses. Wind power profiles $U_{bt}^{\text{WT}}$ and solar power profiles $U_{bt}^{\text{PV}}$ were calculated for each bus $b$ and time $t$ using following equations [28]. These calculations incorporate constants $v^{\text{CI}}$, $v^{\text{R}}$, $v^{\text{CO}}$ and $FF^{\text{PV}}$, which represent the cut-in wind speed, rated wind speed, cut-out wind speed, and fill factor of solar cell, respectively.
@@ II-C KPG 193 test system data
The final KPG 193 test system is illustrated in Fig. 2. The KPG 193 network comprises 193 buses, 123 generators, and 407 transmission lines across four voltage levels. Generator mix and capacity by energy source are visualized using a pie chart at each bus. The network data were formatted to comply with the MATPOWER standard [29], while bus-specific demand data, weather data, and renewable energy profiles over time were organized in CSV format. The resulting synthetic KPG 193 test system dataset is available in [30].
@@ III Validation
TABLE III: Generator Parameters by Fuel Type: The parameters are adopted, randomized, and modified from [31] and [32]. The cost parameters ($C_{g}^{(2)},C_{g}^{(1)},C_{g}^{(0)}$) are presented as ranges across generator units for each fuel type.
@@
LNG 52% 100 4 3 53,862 2.4597 – 8.0681 44,441 – 77,471 779,388 – 7,107,426
Coal 40% 66 6 12 12,606 0.1363 – 0.1663 22,308 – 27,178 1,256,064 – 3,027,738
Nuclear 95% 18 8 12 - 2.1912 – 2.9507 4,095 – 7,193 0
Figure 3: (a) Daily electrical demand and (b) Daily operation costs for the KPG 193 test system, based on 2022 demand and renewable energy data.
The feasibility of the KPG 193 test system was evaluated by solving the daily Unit Commitment (UC) problem [33] and hourly DC Optimal Power Flow (DCOPF) simulations [34] for the entire year of 2022. Figure 3 shows the total demand and optimized operation costs from DCOPF. Since UC parameters and generation cost coefficients are not publicly available, we derived UC parameters from [31] and obtained cost coefficients from a prior study [32], using an exchange rate of 1300 KRW/\$. The quadratic generator cost function is:
$c_{gt}=C^{(2)}_{g}p_{gt}^{2}+C_{g}^{(1)}p_{gt}+C_{g}^{(0)}$ (3)
where $p_{gt}$ is the power output of generator $g$ at time $t$, and $C^{(2)}_{g},C^{(1)}_{g}$, and $C^{(0)}_{g}$ are the cost coefficients. Relevant parameters are detailed in Table III. When solving DCOPF and UC problems, renewable energy curtailment was not permitted, and their associated costs were excluded. Under these conditions, the daily UC and hourly DCOPF simulations converged across 2022 dataset.
@@
To verify these features, we conducted an AC Optimal Power Flow (ACOPF) analysis [34] using snapshot data from 13:00 on July 22, 2022, the peak demand day of the summer. As shown in Fig. 4, the circled metropolitan area accounts for approximately 40% of the total power demand, while generators are concentrated along the coastal regions. This geographic distribution results in northward power flows through 345 kV, 765 kV, and HVDC transmission lines. In Fig. 5, regions with lighter colors indicate lower voltage magnitudes. Along the eastern coast, coal and nuclear power plants are densely clustered, transmitting power over distances exceeding 200 km to the metropolitan area. In contrast, the southwestern region with its high concentration of solar power installations, faces emerging overvoltage issues due to its relatively weak transmission network.
@@ V Conclusion
We present a methodology for devloping a synthetic test system using open-source data to advance power grid modeling in Korea. The data collection process included network topology, generator capacities, weather data, and electrical demand. The KPG 193 test system was created by clustering these datasets based on utility companies’ regional offices, ensuring a computationally tractable model while preserving the key characteristics of the Korean power grid. The system organizes spatial and temporal data from 2022 to provide a comprehensive simulation environment. The feasibility of the test system was validated by solving the daily UC and hourly DCOPF over an entire year. Additionally, a snapshot ACOPF analysis was performed to confirm that the KPG 193 test system replicates the characteristics of the Korean power grid. The simulation results successfully captured power flow and voltage patterns observed in the actual grid.

### v2

@@ Abstract
This paper introduces the 193-bus synthetic Korean power grid (KPG 193), developed using open data sources to address recent challenges of the Korean power system. The KPG 193 test system serves as a valuable platform for decarbonization research, capturing Korea’s low renewable energy penetration, concentrated urban energy demand, and isolated grid structure. Clustering techniques were applied to preserve key system characteristics while maintaining computational tractability and representativeness. The system includes 193 buses, 122 generators, 407 transmission lines, and incorporates temporal weather datasets. Its feasibility was validated through Unit Commitment (UC) and AC Optimal Power Flow (ACOPF) simulations using 2022 demand and renewable generation data. This test system aims to provide a foundational framework for modeling and analyzing the Korean power grid.
Index Terms:
Decarbonization, meteorological dataset, open systems, power system simulation, transmission grid
@@ I Introduction
As of 2022, the Asia-Pacific region accounts for over 50% of global greenhouse gas emissions, with more than 80% of the projected global coal demand growth originating from this region [1]. In Korea, coal contributes 33% of the country’s electricity supply, while renewables account for less than 10% [2]. Given the region’s significant impact on global emissions, advancing open-source power grid modeling efforts in Korea is crucial to support decarbonization studies. Europe has made substantial progress in open-source power grid modeling. The SciGrid [3] and osmTGmod [4] projects modeled Germany’s transmission network using OpenStreetMap (OSM), a publicly available dataset [5]. Building on this, PyPSA-Eur [6] extended the scope to create an open model covering the entire ENTSO-E area. A key enabler in Europe is the public availability of geographical data for substations and transmission lines provided by ENTSO-E [7]. This openness allows for the validation of power grid models built on OSM data, enhancing their accuracy and reliability. In the United States, open test systems have also achieved significant development. The pioneering work at Texas A&M University [8] laid the foundation, later updated the system to meet 2016 standards with additional generators and high-voltage direct current (HVDC) transmission lines [9]. More recently, the TX-123BT network, which represents the Texas power grid and incorporates five years of meteorological data, has been made publicly available [10].
@@
While regions such as Europe and the United States have made significant advancements in open-source power grid modeling through extensive research collaboration, synthetic power grid systems remain underdeveloped in Korea. Previous studies have utilized a realistic 1,428-bus system [11] and a reduced 11-zone system [12]; however, only the 11-zone system is publicly available [13]. To address this gap, we present the synthetic Korean power grid 193-bus test system (KPG 193), developed using open data from the year 2022. Detailed information about the Korean power grid is not publicly disclosed and publicly accessible data is limited to electricity demand and renewable energy capacity at the municipal level. To overcome these limitations while ensuring computational tractability and accurate grid representation, we developed the test system by clustering regions based on the locations of regional offices of the utility company. The proposed KPG 193 test system captures recent challenges of Korea’s power grid, including concentrated electricity demand in metropolitan areas and the geographical distribution of power plants along the coastline. Moreover, the test system can be utilized for decarbonization studies and expansion planning research for Korea. The main contributions of this paper are threefold:
@@
  • Development of the KPG 193 test system: We provide synthetic transmission grid model for Korea based on open-source data, featuring 193 buses, 122 generators, and 407 transmission lines. The resulting KPG 193 is publicly available in [14].
@@
  • Integration of renewables and weather datasets: The system includes weather datasets used to generate renewable generation profile, enabling accurate simulations and facilitating advanced energy system studies.
@@ II-A1 Network Topology
Two types of line data were extracted from the OSM database using the Overpass API: power line for overhead lines and power cable for underground cables [18]. Each line element includes attributes such as voltage level and the number of parallel lines. AC line parameters were calculated using standard conductor coefficients from manufacturer catalogs, considering voltage level and conductor type [19, 20]. A summary of these line parameters is provided in Table II. The Korean power grid also features a HVDC transmission line connecting the Bukdangjin and Godeok substations with a capacity of 1,500 MVA [21]. Since this HVDC line is not included in the OSM database, it was manually added during data preparation.
@@
Overhead line 154 kV ACSR 410 mm×$\times 2$ 2 0.037 0.337 12.9 452
345 kV ACSR 480 mm×$\times 2$ 2 0.031 0.332 13.1 1,086
@@
Capacity [GW] 38.13 41.20 24.65 23.75 7.20 1.65 136.57
Capacity Share [%] 27.9% 30.2% 18.0% 17.4% 5.3% 1.2% 100%
@@ II-B3 Renewable Energy Profile Generation
Renewable energy profiles were generated for each energy source using different methods. The hydropower profile $U_{t}$ was derived from historical generation data provided by KPX [28], representing uniform hydroelectric production across all buses. Wind power profiles $U_{bt}^{\text{WT}}$ and solar power profiles $U_{bt}^{\text{PV}}$ were calculated for each bus $b$ and time $t$ using following equations [29].
@@
These calculations incorporate constants $v^{\text{CI}}$, $v^{\text{R}}$, $v^{\text{CO}}$ and $FF^{\text{PV}}$, which represent the cut-in wind speed, rated wind speed, cut-out wind speed, and fill factor of solar cell, respectively.
@@ II-C KPG 193 test system data
The final KPG 193 test system is illustrated in Fig. 2. The KPG 193 network comprises 193 buses, 122 generators, and 407 transmission lines across four voltage levels. Generator mix and capacity by energy source are visualized using a pie chart at each bus. The network data were formatted to comply with the MATPOWER standard [30], while bus-specific demand data, weather data, and renewable energy profiles over time were organized in CSV format.
@@
TABLE III: Comparison of Annual Electric Generation Share by Fuel Type in 2022: Historical Data and KPG193
Coal Gas Nuclear Renewables Etc. Total
Historical [%] 32.5% 27.5% 29.6% 9.6% 0.8% 100%
KPG193 [%] 33.9% 27.8% 28.2% 10.1% 0% 100%
Etc. includes electricity generated from waste and oil.
@@ III Validation
TABLE IV: Generator Parameters by Fuel Type: The parameters are adopted, randomized, and modified from [31], [32] and [9]. The cost parameters ($C_{g}^{(2)},C_{g}^{(1)},C_{g}^{(0)}$) are presented as ranges across generator units for each fuel type.
@@
LNG 52% 100 4 3 53,862 2.1215 – 6.6711 36,872 – 70,956 637,657 – 6,531,339
Coal 40% 66 6 12 12,606 25.6102 – 30.5675 22,912 – 27,174 1,227,022 – 2,629,634
Nuclear 95% 18 8 12 - 1.6591 – 3.0364 3,339 – 8,292 0
The feasibility of the KPG 193 test system was evaluated by solving the daily Unit Commitment (UC) problem [33] and hourly AC Optimal Power Flow (ACOPF) simulations [34] for the entire year of 2022. Table III presents the validation of generation results through comparison with actual system data [28], with deviations within 1.4%. Since UC parameters and generation cost coefficients are not available, we derived parameters from [31] and [32]. In addition, we modified the cost coefficients using methodologies from [9] to reduce discrepancies with historical data. The quadratic generator cost function is:
@@
$c_{gt}=C^{(2)}_{g}p_{gt}^{2}+C_{g}^{(1)}p_{gt}+C_{g}^{(0)}$ (3)
where $p_{gt}$ is the power output of generator $g$ at time $t$, and $C^{(2)}_{g},C^{(1)}_{g}$, and $C^{(0)}_{g}$ are the cost coefficients. Relevant parameters are detailed in Table IV. When solving UC and ACOPF problems, nuclear generator maintenance data from 2022 is incorporated to reflect actual system conditions [35]. Under these conditions, the daily UC and hourly ACOPF simulations converged across 2022 dataset. The nuclear generator maintenance schedule in the KPG193 is available in [14].
@@
To verify these features, we conducted an ACOPF analysis using snapshot data from 13:00 on July 22, 2022, the peak demand day of the summer. As shown in Fig. 3, the circled metropolitan area accounts for approximately 40% of the total power demand, while generators are concentrated along the coastal regions. This geographic distribution results in northward power flows through 345 kV, 765 kV, and HVDC transmission lines. In Fig. 4, regions with lighter colors indicate lower voltage magnitudes. Along the eastern coast, coal and nuclear power plants are densely clustered, transmitting power over distances exceeding 200 km to the metropolitan area. In contrast, the southwestern region with its high concentration of solar power installations, faces emerging overvoltage issues due to its relatively weak transmission network.
@@ V Conclusion
We present a methodology for devloping a synthetic test system using open-source data to advance power grid modeling in Korea. The data collection process included network topology, generator capacities, weather data, and electrical demand. The KPG 193 test system was created by clustering these datasets based on utility companies’ regional offices, ensuring a computationally tractable model while preserving the key characteristics of the Korean power grid. The system organizes spatial and temporal data from 2022 to provide a comprehensive simulation environment. The feasibility of the test system was validated by solving the daily UC and hourly ACOPF over an entire year. Additionally, a snapshot ACOPF analysis was performed to confirm that the KPG 193 test system replicates the characteristics of the Korean power grid. The simulation results successfully captured power flow and voltage patterns observed in the actual grid.